Grade 11Physical Chemistry

500 ml of 1.2 molar h2so4 is mixed with 1000 ml of 2.5molar solution of h2so4 and the final solution is diluted to 5l . find the molarity of final solution

Shashanka Banerjee
7 Years agoGrade 11

1 Answer

Pooja
7 Years ago
M1V1 + M2V2=MV
(1.2) × 0.5 + (2.5) × 1 =M×5
[Volume is converted to litre ]
0.6 + 2.5 = 5M
Molarity= 0.62M